OpenStudy (anonymous):

A plane flew 1740 km in 6 hours with a tail wind. On the return trip, going into the wind, the plane flew 570 km in 3 hours. What is the plane's speed if there was no wind? What is the speed of the wind? (r+w)6=1740 (r-w)3=570 (r-w)6=1740 (r+w)3=570 (r+w)3= 1740 (r-w)6 = 570 (r-w)3=1740 (r+w)6=570

OpenStudy (goformit100):

(r-w)6=1740 (r+w)3=570

OpenStudy (dariyana):

Let the value v denote the speed of the plane with no wind distance = speed * time We know that 1740 = (v + x) * 6, or (v + x) = 1740 / 6 = 290 Assuming it is the same wind, we also know that 570 = (v - x) * 3, or (v - x) = 570 / 3 = 190 Add the two together and we have: (v + x) + (v - x) = 290 + 190 2v = 480 v = 240 kilometres per hour x = 290 - v = 290 - 240 = 50 So the speed of the plane is 240 km per hour, and the wind is 50 km per hour

OpenStudy (anonymous):

thank you

OpenStudy (dariyana):

you welcome @rebecca1233